Calculating the Predetermined Overhead Rate, Applying Overhead to Production

At the beginning of the year, Debion Company estimated the following:

Overhead $507,500
Direct labor hours 90,000

Debion uses normal costing and applies overhead on the basis of direct labor hours. For the month of March, direct labor hours were 8,300.

This information has been collected in the Microsoft Excel Online file. Open the spreadsheet, perform the required analysis, and input your answers in the questions below.

 

Retrieving data. Wait a few seconds and try to cut or copy again.

Required:

1. Calculate the predetermined overhead rate for Debion. Round your answer to the nearest cent.
$ per direct labor hour
2. Calculate the overhead applied to production in March. Do not round intermediate calculations. Round your answer to the nearest dollar.
$
Answer #1

a. Predetermined overhead rate =Estimated overhead/no of direct labour =507500/90000 =5.64 b. Overhead applied =8300*5.64 =46